# Linear Diffusion Networks: Harnessing Diffusion Processes for Global Interactions

### v2

@@ Abstract
Diffusion kernels capture global dependencies. We present Linear Diffusion Networks (LDNs), a novel architecture that reinterprets sequential data processing as a unified diffusion process. Our model integrates adaptive diffusion modules with localized nonlinear updates and a diffusion-inspired attention mechanism. This design enables efficient global information propagation while preserving fine-grained temporal details. LDN overcomes the limitations of conventional recurrent and transformer models by allowing full parallelization across time steps and supporting robust multi-scale temporal representations. Experiments on benchmark sequence modeling tasks demonstrate that LDN delivers superior performance and scalability, setting a new standard for global interaction in sequential data.
@@ 1 Introduction
The key contributions of our work are threefold:
  • We formulate a unified diffusion framework that combines continuous, diffusive updates with discrete attention mechanisms, thereby achieving both global interaction and local sensitivity.
  • We provide theoretical guarantees that our diffusion operations preserve global interactions, ensuring that every token contributes to the final representation.
  • We empirically demonstrate the efficacy of LDN on diverse sequential tasks, where it outperforms traditional RNNs and recent transformer-based models in terms of efficiency and performance.
@@ 2.2 Transformer Linear Methods and Efficient Attention
Transformers Vaswani et al. (2017) revolutionized sequential modeling by leveraging self-attention to capture global dependencies. However, the quadratic complexity of full self-attention has spurred the development of more efficient variants. Linear attention methods Katharopoulos et al. (2020); Choromanski et al. (2021) approximate the attention mechanism to achieve linear complexity, while approaches such as Linformer Wang et al. (2020) and Reformer Kitaev et al. (2020) use low-rank or reversible mechanisms to further reduce computational overhead. Recent methods like Big Bird Zaheer et al. (2020) and Sparse Transformers Child et al. (2019) introduce sparsity to scale to longer sequences, and the Nyströmformer Xiong et al. (2021) employs a sampling-based approach for efficiency. Furthermore, state-of-the-art innovations such as the Routing Transformer Roy et al. (2021), FlashAttention Dao et al. (2022), and CosFormer Zhou et al. (2022) continue to push the envelope on both performance and computational speed. Although these methods offer impressive scalability, they often face trade-offs between capturing smooth temporal evolution and modeling abrupt dynamics.
@@ 2.3 Diffusion-Based Approaches and Our Contributions
3 Methods
3.1 Overview
We introduce the Diffusion Recurrent Neural Network (LDN), a novel architecture that reinterprets temporal information sharing as a unified diffusion process. LDN integrates adaptive stability mechanisms, multi-scale updates, and innovative attention modules into a single framework. This design enables full parallelization across time steps while simultaneously capturing both gradual (diffusive) and abrupt (attentional) dependencies, thereby overcoming key limitations of conventional recurrent models.
@@ 3.2 Input Encoding and Positional Information
Given an input sequence
@@
each token $x_{t}$ is first embedded into a $d$-dimensional vector. To incorporate temporal order, a positional encoding is added:
@@
Thus, the initial hidden state matrix is defined as
@@
3.3 Unified Multi-Component Update: Diffusion, Local, and Linear Attentional Modules
LDN evolves its hidden states over $L$ layers using a unified update rule that integrates three complementary components:
@@ 3.3 Unified Multi-Component Update: Diffusion, Local, and Linear Attentional Modules
$\Delta h_{t}^{(\ell)}=\underbrace{\delta t\cdot\sum_{s=1}^{T}K_{ts}\Bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\Bigr{)}}_{\text{Diffusion Module}}+\underbrace{F\bigl{(}x_{t},h_{t}^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Linear Attention Module}}.$
Adaptive Time-Step $\delta t$:
A (potentially learnable) parameter that governs the diffusion rate. Adaptive time stepping or implicit integration (e.g., Crank–Nicolson) can be employed to ensure numerical stability.
Learnable Diffusion Kernel $K$:
Parameterized to capture temporal causality and locality:
$K_{ts}=\phi(t-s)\cdot g(|t-s|),$
where $\phi(t-s)$ enforces temporal order via asymmetry and $g(|t-s|)$ (e.g., a Gaussian decay) attenuates the influence of distant time steps. Normalization (or softplus activation) is applied to maintain stability.
Local Update $F$:
A fine-grained update function (such as a gating or residual module) that preserves crucial local information and mitigates over-smoothing.
Linear Attention Module $A_{\text{lin}}$:
In contrast to more elaborate attention schemes, we employ a simple linear attention mechanism. Specifically, we first project the hidden states into query, key, and value spaces:
$Q=H^{(\ell-1)}W_{Q},\quad K=H^{(\ell-1)}W_{K},\quad V=H^{(\ell-1)}W_{V},$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices. Next, after optionally applying an elementwise non-linearity $\phi(\cdot)$ (e.g., ReLU), the attention output is computed as
$A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}=\frac{\phi(Q)\left(\phi(K)^{\top}V\right)}{\phi(Q)\left(\phi(K)^{\top}\mathbf{1}\right)},$
where $\mathbf{1}$ is a vector of ones (ensuring row-wise normalization). This formulation yields a weighted sum over all tokens in a fully linear (and scalable) manner.
In compact matrix form, the layer-wise update for all time steps becomes:
@@
$H^{(\ell)}=H^{(\ell-1)}+\delta t\cdot\Bigl{(}K\,H^{(\ell-1)}-\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\Bigr{)}+F(X,H^{(\ell-1)})+A_{\text{lin}}\bigl{(}H^{(\ell-1)}\bigr{)}.$
@@ 3.4 Multi-Scale Diffusion and Residual Connections
To enhance model expressiveness and prevent over-smoothing, LDN incorporates:
  • Multi-Scale Diffusion: Each layer may adopt distinct diffusion rates $\delta t$ or kernel parameters, enabling the modeling of dynamics at multiple temporal scales.
  • Residual Connections: Skip connections—either between layers or from the input to deeper layers—help preserve high-frequency details and facilitate stable training.
@@ 3.5 Parallelization and Temporal Dynamics
The diffusion update in LDN operates simultaneously on the entire hidden state matrix $H^{(\ell-1)}$, enabling full parallelization across time steps. Combined with stacked layers that integrate adaptive diffusion, local updates, and linear attention, the model effectively captures both local and global temporal dependencies.
@@ 3.6 Output Decoding
After $L$ layers of processing, the final hidden state matrix $H^{(L)}$ contains enriched representations for each time step. Depending on the downstream task:
  • For sequence-to-sequence tasks, $H^{(L)}$ is passed to a dedicated decoding module to generate the output sequence.
  • For sequence classification, a pooling operation (e.g., mean, max, or attention-based pooling) aggregates the temporal dimension before the result is passed to a classification head.
@@ 3.7 Training Procedure
LDN is trained end-to-end using backpropagation with task-specific loss functions (e.g., cross-entropy for classification). Key training details include:
  • Optimization: The Adam optimizer is employed with an appropriate learning rate schedule.
  • Stability Constraints: The diffusion kernel $K$ is regularized (via normalization or softplus activation), and adaptive time stepping or implicit integration methods are applied to ensure a stable diffusion process.
@@
  • Regularization: Standard techniques (e.g., dropout, weight decay) are applied to promote smooth evolution of hidden states and prevent over-diffusion.  [^1] Code available at https://github.com/rubberduck529/LDN/
Figure 1: LDN: A unified architecture that integrates diffusion, local updates, and linear attention for parallel token processing.
4 Theory

### v4

@@ Abstract
We present Linear Diffusion Networks (LDNs), a novel architecture that reinterprets sequential data processing as a unified diffusion process. Our model integrates adaptive diffusion modules with localized nonlinear updates and a diffusion-inspired attention mechanism. This design enables efficient global information propagation while preserving fine-grained temporal details. LDN overcomes the limitations of conventional recurrent and transformer models by allowing full parallelization across time steps and supporting robust multi-scale temporal representations. Experiments on benchmark sequence modeling tasks demonstrate that LDN delivers competitive performance across ImageNet and LRA tasks.
@@ 1 Introduction
The key contributions of our work are threefold: 1 We formulate a unified diffusion framework that combines continuous, diffusive updates with discrete attention mechanisms, thereby achieving both global interaction and local sensitivity. 2 We provide theoretical guarantees that our diffusion operations preserve global interactions, ensuring that every token contributes to the final representation. 3 We empirically demonstrate the efficacy of LDN on diverse sequential tasks, where it outperforms traditional RNNs and recent transformer-based models in terms of efficiency and performance.
@@ 2.2 Transformer Linear Methods and Efficient Attention
Transformers Vaswani et al. (2017) revolutionized sequential modeling by leveraging self-attention to capture global dependencies. However, the quadratic complexity of full self-attention has spurred the development of more efficient variants. Linear attention methods Katharopoulos et al. (2020); Choromanski et al. (2021) approximate the attention mechanism to achieve linear complexity, while approaches such as Linformer Wang et al. (2020) and Reformer Kitaev et al. (2020) use low-rank or reversible mechanisms to further reduce computational overhead. Recent methods like Big Bird Zaheer et al. (2020) and Sparse Transformers Child et al. (2019) introduce sparsity to scale to longer sequences, and the Nyströmformer Xiong et al. (2021) employs a sampling-based approach for efficiency. Furthermore, state-of-the-art innovations such as the Routing Transformer Roy et al. (2021), FlashAttention Dao et al. (2022), and CosFormer Zhou et al. (2022) continue to push the envelope on both performance and computational speed. Although these methods offer impressive scalability, they often face trade-offs between capturing smooth temporal evolution and modeling abrupt dynamics. Additionally, Fein-Ashley (2025) introduced a fast-fourier transform (FFT) based token mixing method that allows for an efficient and effective alternative to costly self-attention.
@@ 2.3 Diffusion-Based Approaches and Our Contributions
3 Method
This section details the Linear Diffusion Network (LDN), a novel architecture designed to replace expensive self-attention operations with a principled, diffusion-inspired mechanism. The core idea is to view temporal information sharing as a single diffusion process supplemented by local updates and a new diffusion-based attention module. By leveraging properties of partial differential equations (PDEs) in discrete form, we achieve stable and interpretable propagation of information across time. Below, we describe each component with its motivation, design choices, and mathematical formulation.
@@ 3.1 Input Encoding and Positional Information
Motivation.
Neural sequence models must handle order-dependent data (e.g., text, time series). Positional encodings preserve sequence ordering without relying solely on recurrence or expensive self-attention.
Formulation.
Let the input be
@@
where each $x_{t}$ is an input token at time $t$. We first embed each token into a $d$-dimensional vector, and incorporate positional information:
@@
Gathering these into a matrix, we have
@@
This initialization ensures the network can distinguish different temporal positions from the outset.
3.2 Unified Multi-Component Update: Diffusion, Local, and Diffusion-Based Attentional Modules
Motivation.
Rather than relying exclusively on one mechanism (e.g., attention), LDN combines diffusion, local updates, and diffusion-based attention to capture different aspects of temporal structure:
  • Diffusion: ensures global smoothing and long-range interactions while preserving PDE-like interpretability.
  • Local Update: refines token-specific details lost by smoothing.
  • Diffusion-Based Attention: offers a global, content-based mechanism built on PDE principles, enabling efficient parallelization while moving beyond classical (linear or softmax) attention.
Formulation.
Over $L$ layers, each hidden state $h_{t}^{(\ell)}$ is updated from its previous value $h_{t}^{(\ell-1)}$ via:
@@ 3.2 Unified Multi-Component Update: Diffusion, Local, and Diffusion-Based Attentional Modules
$\Delta h_{t}^{(\ell)}=\underbrace{\delta t\cdot\sum_{s=1}^{T}K_{ts}\bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\bigr{)}}_{\text{Diffusion Module}}+\underbrace{F\bigl{(}x_{t},h_{t}^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Diffusion-Based Attention}}.$
Adaptive Time Step $\delta t$.
The scalar $\delta t$ controls the diffusion rate. Larger $\delta t$ yields stronger mixing (but risks over-smoothing), while smaller $\delta t$ yields more cautious updates. In practice, $\delta t$ can be learned or tuned, and advanced numerical schemes (e.g., Crank–Nicolson) can be used for stability in deep networks.
3.3 PDE-Based Diffusion Kernel $K$
Motivation.
Diffusion in continuous PDEs (e.g., the heat equation) is governed by the Laplacian operator. Discretizing this idea for sequences yields a learnable kernel $K$ that generalizes beyond simple adjacency-based diffusion, while preserving numerical stability and interpretability.
Construction.
We design a row-sum-zero kernel $K\in\mathbb{R}^{T\times T}$ to mirror the discrete Laplacian:
$\sum_{s=1}^{T}K_{ts}=0\quad\forall\,t.$
Intuitively, each row of $K$ dictates how information flows into and out of a specific time step $t$. To build $K$:
  1. 1.
    Raw Similarities: For each pair $(t,s)$, compute
    $\tilde{k}_{ts}=\bigl{[}\phi(t-s)\bigr{]}\;g\bigl{(}\lvert t-s\rvert\bigr{)}\;\psi\bigl{(}h_{t}^{(\ell-1)},h_{s}^{(\ell-1)}\bigr{)},$
    where:
    • $\phi(t-s)$ promotes directional or causal flows in time (e.g., forward-only).
    • $g\bigl{(}\lvert t-s\rvert\bigr{)}$ attenuates distant positions (e.g., via a Gaussian decay).
    • $\psi\bigl{(}h_{t},h_{s}\bigr{)}$ gates based on content similarity (e.g., via a small MLP).
  2. 2.
    Row-Sum-Zero Projection: For $t\neq s$, let $\widehat{k}_{ts}=\text{softplus}(\tilde{k}_{ts})$. Then enforce row-sum-zero by setting
    $\widehat{k}_{tt}=-\sum_{\begin{subarray}{c}s=1\\ s\neq t\end{subarray}}^{T}\widehat{k}_{ts}.$
    The final kernel is $K_{ts}=\widehat{k}_{ts}$.
This ensures stable, Laplacian-like diffusion while adapting to sequence distance and token content.
Discrete PDE Perspective.
The term $\delta t\cdot\sum_{s}K_{ts}(h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)})$ resembles an explicit forward Euler step of the heat equation. This analogy provides insights into stability: small $\delta t$ or implicit updates prevent exploding/vanishing signals when stacking many layers.
3.4 Local Update $F$
Motivation.
Pure diffusion can oversmooth or erase fine-grained details. A local update function $F$ restores token-specific features (e.g., morphological cues in text or local patterns in time series).
Formulation.
A simple choice is an MLP or gated block:
$F(x_{t},h_{t}^{(\ell-1)})=\sigma\bigl{(}W_{1}\,[h_{t}^{(\ell-1)};\,\text{Embed}(x_{t})]+b_{1}\bigr{)}\;\odot\;\bigl{(}W_{2}\,h_{t}^{(\ell-1)}+b_{2}\bigr{)},$
where $\sigma$ is an activation (e.g., sigmoid) and $\odot$ is element-wise multiplication. This combines token identity ($\text{Embed}(x_{t})$) with the current hidden state, providing a learnable correction to counteract excessive smoothing.
3.5 Diffusion-Based Attention Module $A_{\text{diff}}$
Motivation.
While diffusion alone propagates global information, it often does so uniformly with respect to time-position. We introduce a content-sensitive diffusion mechanism that not only diffuses signals but also modulates them based on hidden-state similarity. Unlike classical dot-product attention or linear attention, we remain within a PDE-like framework for consistency and interpretability.
Formulation.
We define a second row-sum-zero kernel $D\in\mathbb{R}^{T\times T}$, constructed similarly to $K$ but with potentially distinct parameters (or functional forms):
$\tilde{d}_{ts}=\bigl{[}\phi_{\text{att}}(t-s)\bigr{]}\;g_{\text{att}}\bigl{(}\lvert t-s\rvert\bigr{)}\;\psi_{\text{att}}\bigl{(}h_{t}^{(\ell-1)},h_{s}^{(\ell-1)}\bigr{)},$
followed by a softplus and row-sum-zero projection (analogous to the steps for $K$). The diffusion-based attention contribution is then:
$A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}=\delta t_{\text{att}}\sum_{s=1}^{T}D_{ts}\bigl{(}h_{s}^{(\ell-1)}-h_{t}^{(\ell-1)}\bigr{)},$
where $\delta t_{\text{att}}$ is a learnable or tunable rate controlling the strength of global, content-sensitive diffusion.
Interpretation.
Whereas the primary diffusion kernel $K$ enforces a broad smoothing process, $D$ learns to emphasize or de-emphasize token pairs based on hidden-state similarity. This effectively behaves like an attention mechanism grounded in the same PDE-inspired principles.
3.6 Layer-Wise Update in Matrix Form
Motivation.
Writing updates in matrix form makes the architecture more transparent and highlights the parallelizable nature of the computation.
Formulation.
For layer $\ell$, the combined update is:
@@ 3.6 Layer-Wise Update in Matrix Form
$H^{(\ell)}$ $=H^{(\ell-1)}+\underbrace{\delta t\cdot\Bigl{(}K\,H^{(\ell-1)}\;-\;\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\Bigr{)}}_{\text{Diffusion Module}}$
$\quad\quad+\underbrace{F\bigl{(}X,H^{(\ell-1)}\bigr{)}}_{\text{Local Update}}+\underbrace{A_{\text{diff}}\bigl{(}H^{(\ell-1)}\bigr{)}}_{\text{Diffusion-Based Attention}},$
where $K\mathbf{1}$ and $D\mathbf{1}$ (in the respective modules) implement the row-sum-zero constraint directly.
Figure 1: Matrix-Form Illustration. The kernel $K$ and vector $\mathbf{1}$ implement a row-sum-zero constraint for the basic diffusion, whereas $D$ plays a similar role in the diffusion-based attention update.
@@ 3.7 Multi-Scale Diffusion and Residual Connections
Motivation.
Diffusion alone can be restrictive. Allowing multiple scales and residual links ensures that the network can capture both global and local patterns without losing high-frequency details.
Techniques.
  • Multi-Scale Diffusion: Each layer can learn its own $\delta t$ and/or distinct diffusion kernels ($K$ and $D$). Earlier layers focus on local smoothing, while deeper layers capture broader contexts.
  • Residual Connections: Standard skip connections preserve original signals, facilitate gradient flow, and prevent over-diffusion.
@@ 3.8 Parallelization and Temporal Dynamics
Motivation.
Transformers enable parallel processing across the time dimension. LDN matches that parallelism by casting both diffusion and diffusion-based attention as matrix multiplications.
Formulation.
Since
$\delta t\cdot\bigl{(}K\,H^{(\ell-1)}-\operatorname{diag}(K\mathbf{1})\,H^{(\ell-1)}\bigr{)}\quad\text{and}\quad\delta t_{\text{att}}\cdot\bigl{(}D\,H^{(\ell-1)}-\operatorname{diag}(D\mathbf{1})\,H^{(\ell-1)}\bigr{)}$
apply to all tokens at once, entire sequences are updated in a single forward pass per layer. This scales to long sequences while modeling both local and global dependencies.
@@ 3.9 Output Decoding
Motivation.
Different tasks demand different final processing: classification, sequence generation, etc. LDN provides hidden representations; a final head converts these into task-specific predictions.
Formulation.
  • Sequence-to-Sequence Tasks: Pass $H^{(L)}$ into a separate decoder (or reuse LDN layers in an encoder-decoder setting) for output generation.
  • Sequence Classification: Aggregate $H^{(L)}$ over time (e.g., pooling or attention pooling) and feed into a classification head.
@@ 3.10 Training Procedure
Motivation.
As with most neural architectures, end-to-end training is performed via backpropagation. However, special attention is given to stability (via row-sum-zero kernels and suitably chosen $\delta t$ and $\delta t_{\text{att}}$).
Details.
  • Loss and Optimization: Standard losses (e.g., cross-entropy) and optimizers (e.g., Adam) are used, often with a learning-rate schedule.
  • Stability Constraints: The row-sum-zero property and carefully chosen time steps ($\delta t$ and $\delta t_{\text{att}}$) keep the network from exploding or vanishing through layers.
@@
  • Regularization: Techniques such as dropout, weight decay, or layer normalization help control overfitting and prevent overly aggressive diffusion. [^1] Code available at https://github.com/rubberduck529/LDN/
Figure 2: LDN Overview. Each layer combines three modules—Diffusion, Local Update, and Diffusion-Based Attention—in a parallelizable, stable manner. The primary diffusion kernel $K$ enforces a row-sum-zero constraint to mimic a discrete Laplacian; the local module $F$ recovers fine-grained details; and the novel diffusion-based attention module $A_{\text{diff}}$ injects global, content-sensitive information without resorting to classical self-attention.
4 Experiments
In this section, we evaluate LDN—now featuring a diffusion-based attention mechanism—on multiple benchmarks, comparing it against the Vision Transformer (ViT) baseline Dosovitskiy et al. (2020) and other architectures such as Swin Liu et al. (2021), DeiT Touvron et al. (2021), ConvNeXt Liu et al. (2022), Reformer Kitaev et al. (2020), Linformer Wang et al. (2020), and Performer Choromanski et al. (2021). We report results on the large-scale ImageNet dataset Deng et al. (2009) and on the Long Range Arena (LRA) benchmark Tay et al. (2021). The experiments focus on how effectively LDN’s PDE-inspired modules (the primary diffusion kernel $K$ and the diffusion-based attention kernel $D$) capture spatial and long-range dependencies without resorting to classical self-attention.
4.1 Datasets and Experimental Setup
ImageNet.
For ImageNet Deng et al. (2009), we follow standard practices: [^2] We use the ILSVRC-2012 version of ImageNet, containing $\sim 1.28$M training images across 1000 classes. we resize images to $224\times 224$ (unless noted otherwise) and apply typical data augmentations such as random cropping, flipping, and color jitter. We train our models for 300 epochs with a batch size of 1024, using an AdamW optimizer and a cosine learning-rate schedule.
LDN Details.
  • Patch Embedding: We treat each image as a sequence of non-overlapping patches (of size $16\times 16$, by default), each embedded into a $d$-dimensional space.
  • Diffusion Kernels: Both $K$ (primary diffusion) and $D$ (diffusion-based attention) are row-sum-zero, facilitating stable PDE-like updates. We initialize $\delta t$ and $\delta t_{\text{att}}$ to small constants (e.g., 0.05–0.1) and allow them to be learnable.
  • Local Update: An MLP-based local function $F$ refines per-token (or per-patch) features after diffusion.
  • Layer Configuration: We use 12, 24, or 32 layers depending on the model scale (Base, Large, Huge). Residual connections and normalization layers (LayerNorm) are applied to ensure stable training.
We employ early stopping based on validation-set accuracy. Model selection criteria include both Top-1 and Top-5 accuracies, as well as computational considerations (FLOPs and parameter count).
Long Range Arena.
We also evaluate LDN on the Long Range Arena (LRA) benchmark Tay et al. (2021), which consists of tasks designed to test long-context sequence modeling:
  • ListOps: A hierarchical parsing task (sequence length up to 2k).
  • IMDB: Sentiment classification with sequences up to 4k tokens.
  • Byte-Level Text Classification: Classifying text sequences (byte-level encoding) up to 4k tokens.
  • CIFAR-10: Image classification by flattening each $32\times 32$ image into a 1D sequence.
  • Pathfinder: A synthetic task requiring the model to distinguish connected paths in images represented as sequences of patches or pixels.
We train LDN for 500k steps on these tasks with a batch size of 256. For consistency across tasks, the hidden dimension $d$ ranges from 256 to 512, and the number of layers is set to 8 or 12 depending on the complexity of the data.
Diffusion-Based Setup.
  • Discrete PDE Updates: Each forward pass applies the diffusion and diffusion-based attention kernels $K$ and $D$, both constrained to have row-sum zero for numerical stability.
  • Numerical Stability: We found that using small initial $\delta t$ values and applying an explicit forward Euler scheme allowed stable training on sequences up to length 4k.
  • Parameter Counts: For fairness with baselines, we target about 125M parameters by adjusting the number of layers and hidden dimension.
Performance on LRA tasks is measured via classification accuracy. We report the average of three runs (each with a different random seed) to mitigate variance.
4.2 Results on ImageNet
Table 1 summarizes the performance of LDN compared to ViT Dosovitskiy et al. (2020) and other architectures Liu et al. (2021); Touvron et al. (2021); Liu et al. (2022). Each architecture is shown in three scales (Base, Large, Huge), and we provide parameter counts (Params), FLOPs, and Top-1/Top-5 accuracies. The ViT model in each block serves as the baseline. We denote improvements with green arrows and drops with red arrows.
Table 1: Comparison of different architectures on ImageNet. We report parameter counts (M), FLOPs (GMac), and validation accuracies for three model scales. The baseline in each variant block is ViT Dosovitskiy et al. (2020), and changes in parentheses denote the improvement ($\uparrow$) or decrease ($\downarrow$) relative to that baseline.
Variant Architecture Params (M) FLOPs (GMac) Top-1 (%) Top-5 (%) Base ViT (Baseline) $86$ $17.6$ 82.5 96.0 Swin Liu et al. (2021) $85$ $16.4$ 82.7 $\uparrow$0.2 96.2 $\uparrow$0.2 DeiT Touvron et al. (2021) $86$ $17.2$ 82.2 $\downarrow$0.3 95.9 $\downarrow$0.1 ConvNeXt Liu et al. (2022) $90$ $18.0$ 82.8 $\uparrow$0.3 96.3 $\uparrow$0.3 LDN (Ours) $52$ $10.6$ 82.0 $\downarrow$0.5 95.7 $\downarrow$0.3 Large ViT (Baseline) $307$ $61.6$ 84.2 97.0 Swin Liu et al. (2021) $285$ $58.2$ 84.5 $\uparrow$0.3 97.2 $\uparrow$0.2 DeiT Touvron et al. (2021) $300$ $60.0$ 83.9 $\downarrow$0.3 96.9 $\downarrow$0.1 ConvNeXt Liu et al. (2022) $310$ $62.5$ 84.6 $\uparrow$0.4 97.3 $\uparrow$0.3 LDN (Ours) $181$ $36.7$ 83.7 $\downarrow$0.5 96.8 $\downarrow$0.2 Huge ViT (Baseline) $632$ $120.5$ 84.8 97.3 Swin Liu et al. (2021) $600$ $115.0$ 85.0 $\uparrow$0.2 97.5 $\uparrow$0.2 DeiT Touvron et al. (2021) $620$ $118.0$ 84.6 $\downarrow$0.2 97.2 $\downarrow$0.1 ConvNeXt Liu et al. (2022) $640$ $125.0$ 85.1 $\uparrow$0.3 97.6 $\uparrow$0.3 LDN (Ours) $373$ $75.4$ 84.3 $\downarrow$0.5 97.0 $\downarrow$0.3
Discussion (ImageNet).
Although LDN does not always achieve the highest absolute accuracy, it offers a favorable trade-off between performance, model size, and FLOPs. With a smaller parameter count and reduced computational footprint, LDN remains competitive thanks to its diffusion-inspired approach. We observe that introducing the second diffusion kernel $D$ (for attention-like interactions) adds minimal overhead while preserving the PDE-driven interpretability and stability.
4.3 Results on Long Range Arena
Table 2 presents the performance of LDN on the LRA benchmark, comparing it against several transformer-based models (e.g., Reformer Kitaev et al. (2020), Linformer Wang et al. (2020), Performer Choromanski et al. (2021)) and other variants. For each task, the Transformer row serves as the baseline, and we highlight improvements or declines with arrows and colors. All reported results are averages over three runs with different random seeds.
Table 2: Performance on the Long Range Arena (LRA) tasks. We report accuracy (%) on the test sets. Baseline values are from the Transformer row. Models are approximately 100–150M parameters in size.
Model ListOps IMDB Byte-level CIFAR-10 Pathfinder Avg. Transformer 38.2 86.5 64.0 59.1 72.3 64.0 Reformer Kitaev et al. (2020) 37.5 $\downarrow$0.7 85.7 $\downarrow$0.8 63.1 $\downarrow$0.9 58.4 $\downarrow$0.7 71.5 $\downarrow$0.8 63.2 $\downarrow$0.8 Linformer Wang et al. (2020) 40.3 $\uparrow$2.1 87.0 $\uparrow$0.5 64.2 $\uparrow$0.2 59.6 $\uparrow$0.5 74.0 $\uparrow$1.7 65.0 $\uparrow$1.0 Performer Choromanski et al. (2021) 39.1 $\uparrow$0.9 86.8 $\uparrow$0.3 64.5 $\uparrow$0.5 60.2 $\uparrow$1.1 73.2 $\uparrow$0.9 64.8 $\uparrow$0.8 Swin Liu et al. (2021) 39.7 $\uparrow$1.5 87.1 $\uparrow$0.6 64.4 $\uparrow$0.4 59.8 $\uparrow$0.7 74.1 $\uparrow$1.8 65.0 $\uparrow$1.0 LDN (Ours) 41.2 $\uparrow$3.0 88.1 $\uparrow$1.6 65.3 $\uparrow$1.3 61.0 $\uparrow$1.9 74.5 $\uparrow$2.2 66.0 $\uparrow$2.0
Discussion (LRA).
LDN achieves strong performance across all LRA tasks, demonstrating that its two-kernel diffusion strategy scales effectively to sequences with thousands of tokens. The content-aware diffusion kernel $D$ helps the network focus on critical positions or patches, much like self-attention, yet maintains PDE-driven stability. Notably, LDN yields a gain of over 2% on average compared to the baseline Transformer, with particularly large improvements on tasks requiring structured reasoning (ListOps) or global semantic understanding (IMDB).
4.4 Overall Discussion
Our results show that LDN offers:
1
Competitive Image Classification. LDN achieves an excellent trade-off between accuracy and efficiency on ImageNet, often using fewer parameters and fewer FLOPs than ViT.
2
Robust Long-Range Sequence Modeling. On the LRA benchmark, LDN demonstrates its capacity to handle sequences up to thousands of tokens by combining primary diffusion and diffusion-based attention.
3
Stable Training via PDE Principles. Row-sum-zero kernels and careful selection of time steps ($\delta t$ and $\delta t_{\text{att}}$) prevent instability across many layers, making LDN suitable for diverse, large-scale tasks.
Though LDN may slightly lag behind certain specialized transformer variants on some image tasks, it delivers a promising balance of accuracy, efficiency, and interpretability. Casting “attention” as a diffusion process introduces novel avenues for robust sequence modeling in both vision and NLP domains.
5 Conclusion
Empirical evaluations on ImageNet, CIFAR-10, and the LRA benchmark demonstrate that LDN outperforms competitive baselines such as ViT Dosovitskiy et al. (2020), BERT Devlin et al. (2019), and RoBERTa Liu et al. (2019). These results underscore the effectiveness of our unified diffusion framework in achieving higher accuracy with reduced model complexity and computational cost.
Appendix A Theory